OpenStudy (anonymous):

What is the approximate volume of a cone with a height of 5 meters and base radius of 8 meters? Use three point one four for pi. two hundred fifty-one point two cubic meters three hundred thirty-four point nine cubic meters six hundred twenty-eight cubic meters six hundred fifty-three point one two cubic meters

OpenStudy (dangerousjesse):

\[\frac{320 \pi}{3}\approx 335.103\]

OpenStudy (anonymous):

see formula for volume of cone is V=(1/3) * pi * r2 * h or 3.14159 *r2 * l * 4/3 where:pi = constant 3.14159...r = radius of baseh = vertical heightl = length of 'slant' edge of c simply plug in values and get result...

OpenStudy (anonymous):

so it will be 334.9m3....

OpenStudy (anonymous):

option B..

OpenStudy (anonymous):

Okay Thanks! :) May you help me with a little more? :/

OpenStudy (anonymous):

why not ...

OpenStudy (anonymous):

What is the surface area of the figure? A right triangular prism is shown. The side lengths of the triangle measure 6 meters, 8 meters and ten meters. The height of the prism is 3 meters. fifty-seven square meters seventy-two square meters one hundred twenty square meters one hundred thirty-eight square meters

OpenStudy (anonymous):

optin C

OpenStudy (anonymous):

What is the surface area of the composite figure? A rectangular prism with a square base pyramid attached to the top. The square base of the rectangular prism has sides measuring 6 millimeters, the prism's height is 4 millimeters. The pyramid has a square base with side lengths of 6 millimeters, and a slant height of 4 millimeters. sixty square millimeters one hundred eighty square millimeters one hundred ninety-two square millimeters two hundred fifty-two square millimeters

OpenStudy (anonymous):

it should be option C

OpenStudy (anonymous):

Cameron is replacing the roof on a gazebo. The roof is in the shape of a cone with an open bottom. The radius is 6 feet and the slant height is ten feet. What is the lateral area of the roof? Use three point one four for pi. ninety-four point two square feet one hundred point four eight square feet one hundred eighty-eight point four square feet three hundred seventy-six point eight square feet

OpenStudy (anonymous):

option A I THINK

OpenStudy (anonymous):

i have to go now...bye...